# Grounding-IQA: Multimodal Language Grounding Model for Image Quality Assessment

### v2

@@ 3.2 Automated Annotation Pipeline
(2) Open-ended Questions (“What/Why/How”): These questions address low-level attributes or related context (e.g., “What distortion?”); cause analysis (e.g., “Why is quality poor?”); perceptual degree (e.g., “How is clarity?”). All answers are inferred from the description and given as short phrases (e.g., “Noise” and “Medium” ).
Meanwhile, we supply the LLM with the names of key objects/regions (with bounding boxes), constraining the QA to relate to relevant entities. We also use keyword detection to filter out any unrelated QA pairs. Finally, we incorporate bounding box information into the generated QA pairs, forming the corresponding $\bm{question}$ and $\bm{answer}$. This is achieved by applying the same process from Stage-4 of the GIQA-DES construction pipeline.
Figure 5: Some instances from the GIQA-160K dataset, involving two subtasks: GIQA-DES and GIQA-VQA.
@@
Table 1: Statistics information of GIQA-160K and GIQA-Bench. DES: GIQA-DES; VQA: GIQA-VQA.
3.3 GIQA-160K
We construct our grounding-IQA dataset, GIQA-160K, utilizing the automated annotation pipeline, from existing public datasets [43, 52]. Figure 5 contains some instances.
Data Source. To build our dataset, we require two types of data: diverse images and their corresponding detailed quality descriptions. Currently, two public datasets, Q-Pathway [43] and DQ-495K [52], meet our requirements. For Q-Pathway, we select in-the-wild images (KonIQ-10K [9], SPAQ [7], LIVE-FB [50], and LIVE-itw [8]) and AI-generated images (AGIQA-3K [16] and ImageRewardDB [47]), along with their professionally human-annotated texts. The total image-text pairs is 53K. For DQ-495K, the descriptive texts with human annotations correspond to the images in KADIS-700K [20]. These are various types of artificial-degraded images, 27K in total.
@@ 3.4 GIQA-Bench
Bench Statistic. The GIQA-Bench includes 100 images of various types and quality, which are not included in GIQA-160K. We create 100 GIQA-DES and 150 GIQA-VQA test samples based on these images. Among the 150 GIQA-VQA data, 90 are of the “Yes/No” questions (“Yes”: 35; “No”: 55), and 60 are “What/Which/How” questions (“What”: 30; “Why”: 18; “How”: 12). Specifically, the descriptions for GIQA-DES are from Q-Pathway and adjusted, with key objects and corresponding bounding boxes manually determined. GIQA-VQA questions are generated by the annotation pipeline and further refined and answered by humans. Each sample is annotated in multiple rounds by at least three experts with relevant expertise in a controlled laboratory environment to ensure data accuracy.
Evaluation Criteria. We evaluate the grounding-IQA capabilities from three perspectives: (1) Description quality; (2) VQA accuracy; and (3) Grounding precision. For all metrics, higher values indicate better performance.
(1) Description Quality. Assess GIQA-DES performance in quality descriptions. We compare the generated description to the ground truth, excluding coordinates for accuracy. We apply some image captioning metrics: BLEU@4. Moreover, we employ the LLM (i.e., Llama3 [6]) to provide a score from 0 to 4 (higher is better), based on the relevance between the description and the ground truth. For clarity, the final score is scaled proportionally from 0 to 100. We denote the score as the LLM-Score.
(2) VQA Accuracy. Evaluate GIQA-VQA performance in image quality VQA. For “Yes/No” questions, accuracy is determined by matching with the word “Yes” or “No”. For “What/Which/How”, we use LLM to calculate accuracy. The LLM scores the model response from 0 to 4 (higher is better) based on the question and correct answer. The score is then normalized to 0~1. For clarity, we denote the accuracy of “Yes/No” as Acc (Y), “What/Which/How” as Acc (W), and overall accuracy as Acc (Total).
@@
Table 2: Ablation study on box optimization in the automated annotation pipeline. We conduct experiments on the GIQA-DES task.
@@
Table 3: Ablation study on multi-task training.
@@
Table 4: Ablation study on different baselines (data compatibility).
4 Experiments
4.1 Experimental Settings
Implementation Details. We conduct experiments on four pre-trained MLLM models: LLaVA-v1.5-7B [22], LLaVA-v1.5-13B [22], LLaVA-v1.6-7B [23], and mPLUG-Owl2-7B [49]. These models involve different versions, sizes, and architectures. The models are fine-tuned on our proposed GIQA-160K dataset using supervised fine-tuning. We evaluate their performance on grounding-IQA using the GIQA-Bench. Details about the training/testing datasets and evaluation criteria are provided in Secs. 3.3 and 3.4.
Training Settings. We adopt cross-entropy loss for full fine-tuning, following previous methods [43, 21, 49]. The optimizer is AdamW [26], with $\beta_{1}$$=$$0.9$ and $\beta_{2}$$=$$0.999$. We apply the cosine decay scheduler with an initial learning rate of 2$\times$10-5, and a warmup ratio of 0.03. The batch size is set to 64, and the epoch is 2. Other hyper-parameters follow the default settings of each model. Experiments are implemented with PyTorch [32] on four Nvidia A100-80G GPUs.
4.2 Ablation Study
We conduct experiments in this section, analyzing data design and data properties. The training settings are consistent with Sec. 4.1. We apply mPLUG-Owl2-7B [49] as the baseline in all experiments (except in Tab. 4).
Figure 6: Box area distribution of GIQA-160K (Raw and Ref) and GIQA-Bench, showing data in 0~0.5 to highlight differences.
Box Optimization. We evaluate box optimization in annotation pipeline, including box refinement (filter and merge) and representation. We compare the models trained on GIQA-DES with (Ref-Box) and without refinement (Raw-Box) in Tab. LABEL:tab:ablation_box_1. The refinement enhances the fine-tuning effect. We also visualize box area distribution in Fig. 6. Refinement reduces the difference between automatically annotated GIQA-160K and human-annotated GIQA-Bench. Besides, more analyses are provided in the supplement.
Meanwhile, we compare discrete (Disc-Coord) and normalized continuous (Norm-Coord) box representations in Tab. LABEL:tab:ablation_box_2. Results indicate that Disc-Coord enhances description quality (BLEU@4 and LLM-Score) and grounding accuracy (Tag-Recall), compared with Norm-Coord.
Multi-Task Training. We conduct an ablation on multi-task (GIQA-DES/VQA) joint training. The results are listed in Tab. 3. We observe that only GIQA-DES (Only-DES) can improve the quality assessment and grounding, while GIQA-VQA improves VQA accuracy, but grounding ability is limited. Moreover, multi-task training (GIQA-160K) enhances performance on both GIQA-DES/VQA. It demonstrates the importance of data diversity.
Data Compatibility. We fine-tune various baselines using the proposed GIQA-160K. The results are provided in Tab. 4. The results indicate that our dataset is compatible with various MLLMs, effectively enhancing the grounding-IQA ability of the model. Furthermore, we provide more detailed comparisons with more methods in Sec. 4.3.
@@ 4.2 Ablation Study
LLaVA-v1.5-7B [22] N/A N/A 2.82 47.00 N/A N/A 0.4444 0.5167 0.4733
LLaVA-v1.5-13B [22] N/A N/A 3.00 49.00 N/A N/A 0.3888 0.5250 0.4433
LLaVA-v1.6-7B [23] N/A N/A 3.04 50.50 N/A N/A 0.4889 0.5333 0.5067
General mPLUG-Owl2-7B [49] N/A N/A 3.62 48.25 N/A N/A 0.5889 0.5250 0.5633
Shikra-7B [4] 0.4506 0.4768 0.40 27.00 0.4126 0.4359 0.5333 0.3917 0.4767
Kosmos-2-1.6B [33] 0.4946 0.3448 2.63 39.25 0.4982 0.4103 0.3889 0.4750 0.4233
Ferret-7B [51] 0.6458 0.6778 3.16 43.75 0.5393 0.5769 0.4111 0.4875 0.4417
Ground GroundingGPT-7B [18] 0.4967 0.5391 1.99 32.50 0.3845 0.5321 0.5444 0.5250 0.5367
DepictQA-Wild-7B [52] N/A N/A 3.34 56.50 N/A N/A 0.4333 0.5458 0.4783
Q-Instruct [43] (LLaVA-v1.5-7B) N/A N/A 22.69 58.25 N/A N/A 0.6444 0.5375 0.6017
Q-Instruct [43] (LLaVA-v1.5-13B) N/A N/A 19.01 57.25 N/A N/A 0.6222 0.5417 0.5900
IQA Q-Instruct [43] (mPLUG-Owl2-7B) N/A N/A 21.46 62.00 N/A N/A 0.6111 0.5375 0.5817
@@
Table 5: Quantitative results on GIQA-Bench. The best and second-best results are colored red and blue.
Figure 7: Visual comparisons on GIQA-Bench. Our proposed grounding-IQA (blue module) enables more fine-grained quality descriptions (left instance) and QA (right instance) with precise position (i.e., bounding box).
@@ 4.3 Results on GIQA-Bench
In GIQA-Bench, we compare four groups of MLLMs with different functionalities, i.e., (1) General models (General): LLaVA-v1.5-7B [22], LLaVA-v1.5-13B [22], LLaVA-v1.6-7B [23], and mPLUG-Owl2-7B [49]; (2) Multimodal referring and grounding models (Ground): Shikra-7B [4], Kosmos-2-1.6B [33], Ferret-7B [51], and GroundingGPT-7B [18]; (3) IQA models (IQA): DepictQA-Wild-7B [52] and Q-Instruct [43] (fine-tuned three base models); and (4) Our methods (Ours): Four general models fine-tuned on GIQA-160K. The detailed test settings and analyses are provided in the supplementary material.
Quantitative Results. We evaluate all models on GIQA-DES and GIQA-VQA tasks from two aspects: quality assessment and referring/grounding ability, as in Tab. 5. General models perform poorly on both tasks, while task-specific models are more effective in their respective domains. Specifically, grounding MLLMs excel in general grounding tasks but underperform on quality-related objects/areas (GIQA-VQA, Tag-Recall). Conversely, IQA models achieve high description quality (GIQA-DES, LLM-Score), but exhibit low accuracy in GIQA-VQA. In contrast, our method outperforms existing MLLMs. Moreover, to further reveal the IQA capability of our method, we evaluate it on traditional score-based IQA benchmarks. Our method also achieves impressive performance. Details are provided in the supplementary material.
Qualitative Results. We provide some visual comparisons in Fig. 7. For GIQA-DES (left instance), the quality descriptions generated by general (mPLUG-Owl2-7B [49]) and grounding (Ferret [51]) MLLMs are unsatisfactory. In contrast, our method describes image quality more properly with coordinates of key objects affecting the quality. Furthermore, in the GIQA-VQA task (right instance), our method produces more accurate responses to image quality VQA involving spatial perception. More qualitative results are provided in the supplementary material.

### v3

@@ 3.2 Automated Annotation Pipeline
(2) Open-ended Questions (“What/Why/How”): These questions address low-level attributes or related context (e.g., “What types of distortion?”); cause analysis (e.g., “Why the image quality is poor?”); perceptual degree (e.g., “How is clarity?”). All answers are inferred from the description and given as short phrases (e.g., “Noise” and “Medium” ).
Meanwhile, we supply the LLM with the names of key objects/regions (with bounding boxes), constraining the QA to relate to relevant entities. We also use keyword detection to filter out any unrelated QA pairs. Finally, we incorporate bounding box information into the generated QA pairs, forming the corresponding $\bm{question}$ and $\bm{answer}$.
Figure 5: Some instances from the GIQA-160K, involving subtasks: GIQA-DES and GIQA-VQA.
3.3 GIQA-160K
We construct our grounding-IQA dataset, GIQA-160K, utilizing the automated annotation pipeline, from existing public datasets (Wu et al., 2024b; You et al., 2024b). Figure 5 shows some instances.
Data Source. To build our dataset, we require two types of data: diverse images and their corresponding detailed quality descriptions. Currently, two public datasets, Q-Pathway (Wu et al., 2024b) and DQ-495K (You et al., 2024b), meet our requirements. For Q-Pathway, we select in-the-wild images (KonIQ-10K (Hosu et al., 2020), SPAQ (Fang et al., 2020), LIVE-FB (Ying et al., 2020), and LIVE-itw (Ghadiyaram & Bovik, 2015)) and AI-generated images (AGIQA-3K (Li et al., 2023) and ImageRewardDB (Xu et al., 2024)), along with their professionally human-annotated texts. The total image-text pairs is 53K. For DQ-495K, 27K artificially degraded images (from KADIS-700K (Lin et al., 2020)) are paired with human-annotated descriptive texts.
Table 1: Statistics information of the proposed datasets. DES: GIQA-DES; VQA: GIQA-VQA.
@@ 3.4 GIQA-Bench
Bench Statistic. The GIQA-Bench includes 100 images of various types and quality, which are not included in GIQA-160K. We create 100 GIQA-DES and 150 GIQA-VQA test samples based on these images. Among the 150 GIQA-VQA data, 90 are of the “Yes/No” questions (“Yes”: 35; “No”: 55), and 60 are “What/Which/How” questions (“What”: 30; “Why”: 18; “How”: 12).
The descriptions for GIQA-DES are from Q-Pathway and adjusted, with key objects and bounding boxes manually determined. GIQA-VQA questions are generated by the annotation pipeline and further refined and answered by humans. Each sample is annotated in multiple rounds by at least three experts with relevant expertise in a controlled laboratory environment to ensure accuracy.
Evaluation Criteria. We evaluate the grounding-IQA capabilities from three perspectives: description quality, VQA accuracy, and grounding precision. For all metrics, higher values are better.
(1) Description Quality. Assess GIQA-DES performance in quality descriptions. We compare the generated description to the ground truth, excluding coordinates. We apply the image captioning metric: BLEU@4. We also employ the LLM (Llama3 (Dubey et al., 2024)) to provide a score from 0 to 4 (higher is better), based on the relevance between the description and the ground truth. For clarity, the final score is scaled proportionally from 0 to 100. We denote the score as the LLM-Score.
(2) VQA Accuracy. Evaluate GIQA-VQA performance in quality VQA. For “Yes/No” questions, accuracy is determined by matching with the word “Yes” or “No”. For “What/Which/How”, we use LLM to calculate accuracy. The LLM scores the model response from 0 to 4 (higher is better) based on the question and correct answer. The score is normalized to 0~1. We denote the accuracy of “Yes/No” as Acc (Y), “What/Which/How” as Acc (W), and overall accuracy as Acc (Total).
@@
4 Experiments
4.1 Experimental Settings
Implementation Details. We conduct experiments on four pre-trained MLLM models: LLaVA-v1.5-7B (Liu et al., 2024a), LLaVA-v1.5-13B (Liu et al., 2024a), LLaVA-v1.6-7B (Liu et al., 2024b), and mPLUG-Owl2-7B (Ye et al., 2024). These models involve different versions, sizes, and architectures. The models are fine-tuned on our proposed GIQA-160K dataset using supervised fine-tuning. We evaluate their performance on grounding-IQA using the GIQA-Bench. Details about the training/testing datasets and evaluation criteria are provided in Secs. 3.3 and 3.4.
Table 2: Ablation study on box optimization (refinement and representation) in the automated annotation pipeline. We conduct experiments on the GIQA-DES task.
@@ 4.1 Experimental Settings
(b) Box representation.
[…]
tableAblation study on multi-task training. The baseline is the pre-trained model, mPLUG-Owl2-7B, without fine-tuning.
@@
Figure 6: Box area distribution of GIQA-160K (Raw and Ref) and GIQA-Bench.
[…]
4.2 Ablation Study
We analyze method design and data properties. The training settings are detailed in Sec. 4.1. We apply mPLUG-Owl2-7B (Ye et al., 2024) as the baseline in all experiments (except in Tab. 3).
Box Optimization. We evaluate box optimization in the annotation pipeline, including the box refinement (IQA filter and box merge) and the coordinate representation. We compare the models trained on GIQA-DES with (Ref-Box) and without refinement (Raw-Box) in Tab. 2a. The refinement enhances the fine-tuning effect. We also visualize box area distribution in Fig. 6. Refinement reduces the difference between automatically annotated GIQA-160K and human-annotated GIQA-Bench. Besides, more analyses are provided in the supplementary material.
Meanwhile, we compare discrete (Disc-Coord) and normalized continuous (Norm-Coord) box representations in Tab. 2b. Results indicate that Disc-Coord enhances description quality (BLEU@4 and LLM-Score) and grounding accuracy (Tag-Recall), compared with Norm-Coord.
Table 3: Ablation study on different baselines.
@@ 4.2 Ablation Study
Multi-Task Training. We conduct an ablation on multi-task (GIQA-DES and GIQA-VQA) joint training. The results are listed in Tab. 6. We observe that only GIQA-DES (Only-DES) can improve the quality assessment and grounding. GIQA-VQA improves VQA accuracy but exhibits limited grounding ability, likely due to reduced contextual information compared to GIQA-DES. Moreover, multi-task training (GIQA-160K) enhances performance on both GIQA-DES and GIQA-VQA. It demonstrates the importance of data diversity.
Data Compatibility. We fine-tune various baselines using the proposed GIQA-160K. The results are provided in Tab. 3. The results indicate that our proposed dataset is compatible with various MLLMs, effectively enhancing the grounding-IQA ability of the model. Furthermore, we provide more detailed comparisons with more methods in Sec. 4.3.
@@
LLaVA-v1.5-7B N/A N/A 2.82 47.00 N/A N/A 0.4444 0.5167 0.4733
LLaVA-v1.5-13B N/A N/A 3.00 49.00 N/A N/A 0.3888 0.5250 0.4433
LLaVA-v1.6-7B N/A N/A 3.04 50.50 N/A N/A 0.4889 0.5333 0.5067
General mPLUG-Owl2-7B N/A N/A 3.62 48.25 N/A N/A 0.5889 0.5250 0.5633
Shikra-7B 0.4506 0.4768 0.40 27.00 0.4126 0.4359 0.5333 0.3917 0.4767
Kosmos-2-1.6B 0.4946 0.3448 2.63 39.25 0.4982 0.4103 0.3889 0.4750 0.4233
Ferret-7B 0.6458 0.6778 3.16 43.75 0.5393 0.5769 0.4111 0.4875 0.4417
Ground GroundingGPT-7B 0.4967 0.5391 1.99 32.50 0.3845 0.5321 0.5444 0.5250 0.5367
DepictQA-Wild-7B N/A N/A 3.34 56.50 N/A N/A 0.4333 0.5458 0.4783
Q-Instruct (LLaVA-v1.5-7B) N/A N/A 22.69 58.25 N/A N/A 0.6444 0.5375 0.6017
Q-Instruct (LLaVA-v1.5-13B) N/A N/A 19.01 57.25 N/A N/A 0.6222 0.5417 0.5900
IQA Q-Instruct (mPLUG-Owl2-7B) N/A N/A 21.46 62.00 N/A N/A 0.6111 0.5375 0.5817
@@
Table 4: Quantitative results on GIQA-Bench. Best and second-best results are colored red and blue.
Figure 7: Visual comparisons on GIQA-Bench. Our proposed grounding-IQA (blue module) enables more fine-grained quality descriptions (left instance) and QA (right instance).
@@ 4.3 Results on GIQA-Bench
In GIQA-Bench, we compare four groups of MLLMs with different functionalities, i.e., (1) General models (General): LLaVA-v1.5-7B (Liu et al., 2024a), LLaVA-v1.5-13B (Liu et al., 2024a), LLaVA-v1.6-7B (Liu et al., 2024b), and mPLUG-Owl2-7B (Ye et al., 2024); (2) Multimodal referring and grounding models (Ground): Shikra-7B (Chen et al., 2023), Kosmos-2-1.6B (Peng et al., 2024), Ferret-7B (You et al., 2024a), and GroundingGPT-7B (Li et al., 2024); (3) IQA models (IQA): DepictQA-Wild-7B (You et al., 2024b) and Q-Instruct (Wu et al., 2024b) (fine-tuned three base models); and (4) Our methods (Ours): Four general models fine-tuned on GIQA-160K. The detailed test settings and analyses are provided in the supplementary material.
Quantitative Results. We evaluate all models on GIQA-DES and GIQA-VQA from two aspects: quality assessment and grounding ability, as in Tab. 4. General models perform poorly on both tasks, while task-specific models are more effective in their respective domains. Specifically, grounding MLLMs excel in grounding tasks but underperform on quality-related objects/areas (GIQA-VQA, Tag-Recall). Conversely, IQA models achieve high description quality (GIQA-DES, LLM-Score), but exhibit low accuracy in GIQA-VQA. In contrast, our method outperforms existing MLLMs.
Moreover, to further demonstrate the performance and generalization ability of our approach, we conduct extensive experiments and evaluations in the supplementary material, including: (1) traditional score-based IQA tasks; (2) the user study on GIQA-Bench, and (3) the application of grounding-IQA to downstream tasks. Our method also achieves impressive performance.
Qualitative Results. We provide some visual comparisons in Fig. 7. For GIQA-DES (left instance), the quality descriptions generated by general (mPLUG-Owl2-7B (Ye et al., 2024)) and grounding (Ferret (You et al., 2024a)) MLLMs are unsatisfactory. In contrast, our method describes image quality more properly with coordinates of key objects affecting the quality. Furthermore, in the GIQA-VQA task (right instance), our method produces more accurate responses to image quality VQA involving spatial perception. More results are provided in the supplementary material.